OpenStudy (anonymous):

Please check my work! Find the Laplace Transform of the given piecewise function:

OpenStudy (anonymous):

\[\huge f(t)=\left\{\begin{matrix} 0, & t < 2\\ (t-2)^2, & t \ge 2 \end{matrix}\right.\] \[\huge \mathcal{L}[(t-2)^2u_2(t)] =e^{-2s} \mathcal{L}[t^2]=2e^{-2s} \cdot \frac{1}{s^3}\]

OpenStudy (anonymous):

@ganeshie8 @IrishBoy123 @freckles X)

OpenStudy (anonymous):

Woops, the end got cut off :( \[\huge 2e^{-2s} \cdot \frac{1}{s^3}\]

OpenStudy (irishboy123):

haven't looked at the table but that seems great @Astrophysics

ganeshie8 (ganeshie8):

\[ f(t)=\left\{\begin{matrix} 0, & t < 2\\ (t-2)^2, & t \ge 2 \end{matrix}\right.\] \[ \mathcal{L}[(t-2)^2u_2(t)] =e^{-2s} \mathcal{L}[t^2]=2e^{-2s} \cdot \frac{1}{s^3}\]

OpenStudy (anonymous):

Okay, I was just checking! I make the dumbest mistakes I tell you.. hehe

OpenStudy (astrophysics):

Learning laplace as well eh :D, me too

OpenStudy (anonymous):

@Astrophysics It's really not fun :( I didn't mind the other sections of differential eq's though!

OpenStudy (astrophysics):

Yeah! I feel exactly the same way!